We have 2 varieties of concrete that we can mix. One is 1 part water to 3 parts concrete mix, and the other is 1 part water to 5 parts concrete mix. We want to end up with 30 lbs of a result that is 1 parts water and 4 part concrete. Determine how many pounds of each variety we need.
Answer by srinivas.g(540)

first variety : portion of water = 1/4 (since ratio = 1:3 total 4 parts)
second variety : portion of water = 1/6 (since ratio = 1:5 total 6 parts)
mix : portion of water = 1/5 (since ratio = 1:4 total 5 parts)

1/4 1/6

1/5
proportion of first variety = (1/5 - 1/6 ) = 1/30
proportion of first variety = (1/4 - 1/5 ) = 1/20
ratio in which first variety and second variety to be mixed = 1/30 : 1/20
=2:3
total volume of the mix = 30 lbs
volume of the first variety int the mix = 2/5 *(30 )
=12 lbs
volume of the second variety int the mix = 3/5 * 30
= 18 lbs